# Unpacking Approaches to Learning and Teaching Machine Learning in K-12 Education: Transparency, Ethics, and Design Activities

### v1

@@ 4.2. Learning algorithm-driven approach
A different approach centers on glassboxing how learning algorithms work and blackboxing the datawork involved in the ML pipeline. We only identified seven different studies that take a learning algorithm-driven approach. In these studies, learners were introduced to neural networks, nearest neighbor, k closest neighbors, NLP methods, and search algorithms, among others, using off-the-shelf datasets (e.g., datasets from Kaggle). These studies were all conducted with secondary school students in (Sperling and Lickerman, 2012) and out of school (Norouzi et al., 2020).
While some studies involved students learning about ML in computing settings such as CS classes and summer camps (Norouzi et al., 2020; Sperling and Lickerman, 2012; Alvarez et al., 2022), others integrated ML into science curriculum (Akram et al., 2022). For students in computing contexts ML learning activities required some pre-existing knowledge of programming. As such, camps included python bootcamps (Norouzi et al., 2020) or introductory python activities (Alvarez et al., 2022) and classroom interventions involving students with more than 2 years of traditional computing education and knowledge of python (Sperling and Lickerman, 2012). On the other hand, more introductory activities in non-computing contexts used block programming environments or no programming at all. Akram et al. (Akram et al., 2022), for instance, presented curricular modules that integrate ML using a block programming environment into early secondary school (grades 6-8) science. Each module involves a science topic, a core ML algorithm, and an algorithmic justice topic. For example one module involves using the bread-first search algorithm, uniform cost search and adversarial search for contact tracing in the context of Covid19 while discussing issues of privacy. Other algorithms include knowledge-based systems, clustering with K-means, decision trees, NLP methods (feature extraction, information retrieval, feature selection, classification). Priya and colleagues (Priya et al., 2022), on the other hand, created a game that provides a conceptual overview of supervised learning, gradient descent, and K-Nearest Neighbor classification without opportunities to build and experiment with models.
@@ 4.2.1. Tools & Curricula
Studies involved curricular activities of different length from 90 minutes (Akram et al., 2022) to 450 hours (Sperling and Lickerman, 2012). Akram and colleagues’ (Akram et al., 2022) ML modules for science classrooms are short by design so that they can be integrated when learning about epidemics, natural disasters, and gravity. Norouzi et al. (Norouzi et al., 2020) presented a four week curriculum that included one week of coding exercises with python, one week on NLP techniques, and one week for working on projects and guest lectures and field trips. Sperling’s (Sperling and Lickerman, 2012) 450 hour course introduced theoretical ideas of algorithms, machine learning, ML algorithms and provided opportunities for students to create full-fledged final projects. Two of the studies we reviewed introduced novel extensions to the Snap programming environment that enable students to use and explore ML algorithms (Akram et al., 2022; Alvarez et al., 2022).
@@ 4.2.2. Making models in the context of designing applications
Studies in the algorithm-driven approach involved students in using different ML algorithms and techniques to analyze data and sometimes create applications. Because most studies use off-the-shelf datasets, many of the projects and applications that students worked on were predetermined by curricular designers and researchers (Akram et al., 2022; Norouzi et al., 2020). These involved having students detect breast cancer or malaria (Norouzi et al., 2020), or build covid contact tracing and disaster risk prevention applications (Akram et al., 2022). Alvarez et al., on the contrary, provided students with open-ended activities in which they could use NLP to analyze any data of their choice from Genius Lyrics, Twitter, and the New York Times APIs. In Sperling (Sperling and Lickerman, 2012) as part of a 450 hours curriculum students created their own application of their choice using an ML algorithm, built a user interface, and examined the performance of their project.
@@ 4.3.1. Mix & match
We identified 15 different studies that integrated both approaches often involving having different hands-on learning activities that focus on either approach. Long and Magerko (Long et al., 2019, 2021a), for example, involved youth in learning about ML with unplugged activities in museum settings. They integrated both approaches by creating distinct exhibit experiences focused on ML competencies. For example, some exhibits centered on explaining how neural networks work and their ways of representing knowledge while others prioritized engaging youth with data and how data shapes system behaviors.
In formal school settings, Lee and colleagues (Lee and Perret, 2022) present an ML methods in data science curriculum in which high school students first engage with a unit focused on data, exploring datasets, considering ethical issues of data production, and analyzing data, build datasets for image classification and later participate in activities designed to learn about perceptrons, neural networks, back propagation, transfer learning, and K nearest networks. Of note data design issues are also integrated in the activities related to learning algorithms as the curriculum focuses on the whole ML pipeline. In a different curriculum Lee and colleagues (Lee et al., 2021b) offer a comprehensive 30 hour introduction to ML for middle schoolers that involves both data-driven units and algorithm-driven units. From creating data sets to train classifiers to exploring the structure of neural networks to an introduction to general adversarial networks. Here every unit also includes an ethics component.
Researchers have also designed tools to support sensemaking of how data and learning algorithms work together in classification tasks. Mahipal et al. (Mahipal et al., 2023) present a curriculum for middle schoolers that integrates both data and algorithmic driven approaches. In this curriculum students use DoodleIt, a tool that supports students to explore and visualize how data is processed by a neural network. This tool, while classifying drawings in real time, visualizes the application of the kernels of a convolutional neural network to the data by creating feature maps.
@@ 4.3.2. Data-driven with algorithm sprinkles
Eight studies involved students in data-driven activities (such as creating data sets to train and then test models) while also explaining the learning algorithms used in such models through lectures and videos. In How To Train Your Robot (Williams et al., 2021), for example, hands-on learning activities involve creating classifiers for images and text which are accompanied by explanations of the k-nearest neighbor algorithm and how it is used. Shamir and Levin (Shamir and Levin, 2022) had students train and test classifiers and later design a logic gate to simulate a (single) artificial neuron. Kaspersen et al. (Kaspersen et al., 2022) introduced VotestratesML, a tool that supports students in processing and creating datasets to train models in the context of social studies classes. This data-driven tool enables learners to choose if they want to train their models using k-nearest neighbors or feedforward neural networks, as well as the parameters used in each algorithm (the value of k, the number of training iterations). As such, students can compare the performance of models trained with both algorithms.
@@ 4.3.3. Algorithm-driven with data sprinkles
Three studies focused on learning algorithms and included discussions and activities about how data influences model behavior. Here for instance, Wan et al. (Wan et al., 2020) and Zhou et al.(Zhou et al., 2021) had students interact with SmileyCluster, a tool that supports students to make sense of the k-means clustering algorithm. Here, while students interacted with data, the goal of the activities was to support learners in understanding the k-means clustering method. In another study, Reddy et al. (Reddy et al., 2021) engaged students with a curriculum that emphasized word embeddings, the k-nearest neighbors algorithm and classification bias. Here students built datasets for the programming activity of their final project, but datasets were not the focus of the intervention.
@@ 5. Discussion
Through our review of the literature we identified and conceptualized three approaches to learning and teaching ML in K-12. Here we discuss the implications of what we observed in each approach. First, we observed that the majority of efforts in ML education have focused on data driven approaches with much less attention given to the role that learning algorithms play in ML model performance. The data driven approach glassboxes the role of training data in shaping model behaviors, current efforts within this approach include learning activities for K-12 students of all ages. On the other hand, all studies in the learning algorithm-driven approach, which glassboxes how machine learning algorithms work, centered on secondary school students with previous experience with programming and more advanced knowledge of mathematics. This disparity highlights the need to develop learning activities and curricula that make learning algorithms accessible to younger students.
@@
Third, we know that, in computing education, creating applications, particularly personally relevant applications, can provide a context that motivates students to learn computing (Waite and Sentance, 2021; Oleson et al., 2020). Yet many efforts in ML education prioritize learning about and creating models in isolation without incorporating them into applications. Almost half of the studies in the data driven approach involve having students build applications, providing opportunities for learning both ML and traditional computing education practices and concepts. In the algorithm-driven approach, while some efforts supported students to use learning algorithms with data sets of their own interest, others involved using datasets selected by researchers and instructors, limiting the opportunities to create personally relevant applications.

### v2

@@ 4.2. Learning algorithm-driven approach
A different approach centers on glassboxing how learning algorithms work and blackboxing the datawork involved in the ML pipeline. We only identified seven different studies that took a learning algorithm-driven approach. In these studies, learners were introduced to neural networks, nearest neighbor, k closest neighbors, NLP methods, and search algorithms, among others, using off-the-shelf datasets (e.g., datasets from Kaggle). These studies were all conducted with secondary school students in (Sperling and Lickerman, 2012) and out of school (Norouzi et al., 2020).
While some studies involved students learning about ML in computing settings such as CS classes and summer camps (Norouzi et al., 2020; Sperling and Lickerman, 2012; Alvarez et al., 2022), others integrated ML into science curriculum (Akram et al., 2022). For students in computing contexts ML learning activities required some pre-existing knowledge of programming. As such, camps included python bootcamps (Norouzi et al., 2020) or introductory python activities (Alvarez et al., 2022) and classroom interventions involved students with more than 2 years of traditional computing education and knowledge of python (Sperling and Lickerman, 2012). On the other hand, introductory activities in non-computing contexts used block programming environments or no programming at all. Akram et al. (Akram et al., 2022), for instance, presented curricular modules that integrate ML using a block programming environment into early secondary school (grades 6-8) science. Each module involved a science topic, a core ML algorithm, and an algorithmic justice topic. For example one module involved using the bread-first search algorithm, uniform cost search and adversarial search for contact tracing in the context of Covid-19 while discussing issues of privacy. Other algorithms included knowledge-based systems, clustering with K-means, decision trees, NLP methods (feature extraction, information retrieval, feature selection, classification). Priya and colleagues (Priya et al., 2022), created a game that provided a conceptual overview of supervised learning, gradient descent, and K-Nearest Neighbor classification without opportunities to build and experiment with models.
@@ 4.2.1. Tools & Curricula
Studies involved curricular activities of different length from 90 minutes (Akram et al., 2022) to 450 hours (Sperling and Lickerman, 2012). Akram and colleagues’ (Akram et al., 2022) ML modules for science classrooms were short by design so that they can be integrated when learning about epidemics, natural disasters, and gravity. Norouzi et al. (Norouzi et al., 2020) presented a four week curriculum that included one week of coding exercises with python, one week on NLP techniques, and one week for working on projects, guest lectures and field trips. Sperling’s (Sperling and Lickerman, 2012) 450 hour course introduced theoretical ideas of algorithms, machine learning, ML algorithms and provided opportunities for students to create full-fledged final projects. Two of the studies we reviewed introduced novel extensions to the Snap programming environment that enabled students to use and explore ML algorithms (Akram et al., 2022; Alvarez et al., 2022).
@@ 4.2.2. Making models in the context of designing applications
Studies in the algorithm-driven approach involved students in using different ML algorithms and techniques to analyze data and sometimes create applications. Because most studies used off-the-shelf datasets, many of the projects and applications that students worked on were predetermined by curricular designers, instructors, and researchers (Akram et al., 2022; Norouzi et al., 2020). These involved, for example, having students detect breast cancer or malaria (Norouzi et al., 2020), or build covid contact tracing and disaster risk prevention applications (Akram et al., 2022). Alvarez et al., on the contrary, provided students with open-ended activities in which they could use NLP to analyze any data of their choice from Genius Lyrics, Twitter, and the New York Times APIs. In Sperling (Sperling and Lickerman, 2012), as part of a 450 hours curriculum students created their own application of their choice using an ML algorithm, built a user interface, and examined the performance of their project.
@@ 4.3.1. Mix & match
We identified 15 different studies that integrated both approaches often having different hands-on learning activities that focus on either approach. Long and Magerko (Long et al., 2019, 2021a), for example, involved youth in learning about ML with unplugged activities in museum settings. They integrated both approaches by creating distinct exhibit experiences focused on ML competencies. For example, some exhibits centered on explaining how neural networks work and their representations while others prioritized engaging youth with data and how data shapes system behaviors.
In formal school settings, Lee and colleagues (Lee and Perret, 2022) presented an ML methods in data science curriculum in which high school students first engaged with a unit focused on data, exploring datasets, considering ethical issues of data production, and analyzing data, building datasets for image classification and later participated in activities designed to learn about perceptrons, neural networks, back propagation, transfer learning, and K nearest networks. Of note data design issues were also integrated in the activities related to learning algorithms as the curriculum focused on the whole ML pipeline. In a different curriculum Lee and colleagues (Lee et al., 2021b) offered a comprehensive 30 hour introduction to ML for middle schoolers that included both data-driven units and algorithm-driven units. From creating data sets to train classifiers to exploring the structure of neural networks to an introduction to generative adversarial networks. Here every unit also included an ethics component.
Researchers have also designed tools to support sensemaking of how data and learning algorithms work together in classification tasks. Mahipal et al. (Mahipal et al., 2023) presented a curriculum for middle schoolers that integrated both data and algorithmic driven approaches. In this curriculum students used DoodleIt, to explore and visualize how data is processed by a neural network. This tool, while classifying drawings in real time, visualizes the application of the kernels of a convolutional neural network to the data by creating feature maps.
@@ 4.3.2. Data-driven with algorithm sprinkles
Eight studies involved students in data-driven activities (such as creating data sets to train and then test models) while also explaining the learning algorithms used in such models through lectures and videos. In How To Train Your Robot (Williams et al., 2021), for example, hands-on learning activities involved creating classifiers for images and text accompanied by explanations of the k-nearest neighbor algorithm and how it is used. Shamir and Levin (Shamir and Levin, 2022) had students train and test classifiers and later design a logic gate to simulate a (single) artificial neuron. Kaspersen et al. (Kaspersen et al., 2022) introduced VotestratesML, a tool that supports students in processing and creating datasets to train models in the context of social studies classes. This data-driven tool enabled learners to choose if they wanted to train their models using k-nearest neighbors or feedforward neural networks, as well as the parameters used in each algorithm (the value of k, the number of training iterations). As such, students could compare the performance of models trained with both algorithms.
@@ 4.3.3. Algorithm-driven with data sprinkles
Three studies focused on learning algorithms and included discussions and activities about how data influences model behavior. Here for instance, Wan et al. (Wan et al., 2020) and Zhou et al.(Zhou et al., 2021) had students interact with SmileyCluster, a tool that supports students to make sense of k-means clustering. Here, while students interacted with data, the goal of the activities was to support learners in understanding the k-means clustering method. In another study, Reddy et al. (Reddy et al., 2021) engaged students with a curriculum that emphasized word embeddings, the k-nearest neighbors algorithm and classification bias. Here students built datasets for the programming activity of their final project, but datasets were not the focus of the intervention.
@@ 5. Discussion
Through our review of the literature we identified and conceptualized three approaches to learning and teaching ML in K-12. Here we discuss the implications of what we observed in each approach. First, we observed that the majority of efforts in ML education have focused on data driven approaches with much less attention given to the role that learning algorithms play in ML model performance. The data driven approach glassboxes the role of training data in shaping model behaviors, current efforts within this approach include learning activities for K-12 students of all ages. On the other hand, all studies in the learning algorithm-driven approach, which glassboxes how learning algorithms work, centered on secondary school students with previous experience with programming and more advanced knowledge of mathematics. This disparity highlights the need to develop learning activities and curricula that make learning algorithms accessible to younger students.
@@
Third, we know that, in computing education, creating applications, particularly personally relevant applications, can provide a context that motivates students to learn computing (Waite and Sentance, 2021; Oleson et al., 2020). Yet many efforts in ML education prioritize learning about and creating models in isolation without incorporating them into applications. Almost half of the studies in the data driven approach involved having students build applications, providing opportunities for learning both ML and traditional computing education practices and concepts. In the algorithm-driven approach, while some efforts supported students to use learning algorithms with data sets of their own interest, others involved using datasets selected by researchers and instructors, limiting the opportunities to create personally relevant applications.